OpenStudy (anonymous):

What is the solution set -1/4x >or equal to -1/8

OpenStudy (anonymous):

multiply each side by -4 \[x <1/2\]

OpenStudy (zarkon):

is the question \[-\frac{1}{4}x\geq -\frac{1}{8}\]?

OpenStudy (anonymous):

yes Zarkon, how did you make that sign?

OpenStudy (zarkon):

\geq

OpenStudy (zarkon):

using the equation editor

OpenStudy (zarkon):

\leq \[\leq\] for the opposite direction

OpenStudy (zarkon):

\[x\leq \frac{1}{2}\]

OpenStudy (anonymous):

yeah I used wrong sign

OpenStudy (anonymous):

Thanks